OpenStudy (anonymous):

Suppose that in one week, your company takes in $2,220 in sales at one of its cell phone stores. The price breakdown per phone is as follows: Cell phone model A4: $35 Smartphone Z20: $50 Suppose that a total of 51 phones were sold. a.Set up the system of equations that needs to be solved to determine how many of each type of phone were sold. Give a clear definition of the variables in the system. b.Solve the system of equations, showing clearly how the solution was determined, and state the results clearly in light of the real-world situation. c.Verify your results of the 2 linear e

OpenStudy (anonymous):

X + Y = 51....1 35X + 50Y = 2200.......2 where X is Cell phone A4 and Y is Z20 Dividing by 5 7X + 10Y = 440....3 7X + 7Y = 357...1*3 - - - 3Y=83 Y=27.67 (Considering it cant be in decimal) X=51-28=23